what are the X and y intercepts of x^2+5x+6

Answer by nerdybill(7384)

to find the y-intercept(s) set x=0 and solve for y:
y=x^2+5x+6
y=0^2+5(0)+6
y=0+0+6
y=6
y-intercept is at (0,6)
.
to find the x-intercept(s) set y=0 and solve for x:
y=x^2+5x+6
0=x^2+5x+6
0=(x+2)(x+3)
x = {-2,-3}
x-intercepts are at (-2,0) and (-3,0)